# Plasma Instabilities Dominate Radioactive Transients Magnetic Fields: The self-confinement of leptons in Type Ia and Core-Collapse Supernovae, and Kilonovae

### v1

@@ 1 Introduction
SNe Ia, in particular, are ideal environments for understanding lepton propagation in radioactively-powered ejecta. Despite the observational necessity, the mechanism behind the continued positron confinement in SNe Ia is unclear. Previous studies have suggested that extremely strong pre-existing magnetic fields ($>10^{6}\,\mathrm{G}$) can explain local positron trapping at late times ($>300$ days; Milne et al., 1999; Penney & Hoeflich, 2014; Hristov et al., 2021), but implies initial WD surface fields of $>10^{9}\,\mathrm{G}$, surpassing all observed field strengths ($10^{3}-10^{9}\,\mathrm{G}$; Schmidt et al., 2003; Külebi et al., 2009; Ferrario et al., 2015). Another option is the delayed deposition of positron energy, which is predicted to flatten the light curve (Kushnir & Waxman, 2020), but has yet to be observationally confirmed.
@@
Plasma streaming instabilities offer a promising solution for positron confinement without the need for strong pre-existing magnetic fields. Plasma streaming instabilities draw energy from the bulk motion of one species and channel it into amplifying electromagnetic perturbations over time, disrupting the plasma’s equilibrium and leading to strong and often turbulent fields. These instabilities play a crucial role in the dynamics of high-energy astrophysical phenomena, but are often overlooked in transient simulations. They are instrumental in processes ranging from electron scattering in the solar wind (Roberg-Clark et al., 2018; Cattell et al., 2021) and intracluster medium of galaxy clusters (Roberg-Clark et al., 2018) to forward shock of $\gamma$-ray bursts (Sironi & Spitkovsky, 2011; Golant et al., 2024). Among the most pertinent plasma streaming instabilities arising from a net electron or positron current are filamentation instabilities (e.g., Weibel, 1959; Bell, 1978, 2004; Bret, 2009; Reville & Bell, 2012; Caprioli & Spitkovsky, 2013; Gupta et al., 2021). Such instabilities can grow exponentially from the density perturbations in the anisotropic flow of particles even without an initial magnetic field and saturate when the electron or positron gyro-radius (Larmor radius) becomes comparable to their inertial length. They can efficiently trap energetic particles by generating local magnetic fields that slow down or halt particle diffusion (Riquelme & Spitkovsky, 2011; Schroer et al., 2021; Gupta et al., 2021).
@@
In this work, we study plasma streaming instabilities in radioactive transients using simulations, taking SNe Ia as fiducial examples. Using fully kinetic particle-in-cell (PIC) simulations, we demonstrate how these instabilities arise and evolve in the presence of high-energy leptons, leading to self-generated magnetic fields that confine these particles even in the absence of an initial magnetic field. In Section 2, we establish the theoretical groundwork of plasma instabilities generated by high-energy leptons. In Section 3, we describe our particle-in-cell (PIC) simulations and the results. In Section 4, we apply these findings to radioactive transients such as SNe Ia, CC SNe, and KNe. Finally, in Section 5, we summarize our conclusions.
@@ 3 Particle-in-Cell Simulations
(Spitkovsky, 2005; Hakobyan et al., 2023), a publicly available, fully parallelized particle-in-cell (PIC) code specifically designed for astrophysical plasma simulations. The PIC method solves the coupled Vlasov-Maxwell system of equations for a collisionless plasma. Maxwell’s equations
@@ 3.1.2 Simulation Results
From this simple simulation with an initial magnetic field of zero, the system generates a saturated magnetic field in microseconds perpendicular perpendicular to the bulk flow. Since positrons are constrained to move along magnetic field lines, the perpendicular fields suggests that the positrons will remain fixed at the same mass coordinates in a radially symmetric explosion. To move outward, they must diffuse, indicating that they would deposit their energy locally instead.
@@
To examine how the particles are affected by the growing magnetic field, we show the initial and final distributions of particles in the two-dimensional velocity-space of $v_{x}$ and $v_{y}$, in Figure 4a and b. Due to the symmetry of the problem, $v_{y}$ and $v_{z}$ have the same statistical properties. The contours show the distributions of positrons, electrons, and ions. The inner levels of the contour represent a higher density of particles in velocity-space.
@@
Figure 4: (a) The initial and (b) final ($t=540\,\mathrm{\omega_{pe}^{-1}}$) velocity distributions for the positrons (blue), electrons (red), and ions (green) for the fiducial SN Ia simulation. The velocity of $v_{x,y}=0$ is marked by the dashed gray lines, indicating the point at which the particles would be at rest. (c) The initial and (d) final ($t=540\,\mathrm{\omega_{pe}^{-1}}$) kinetic energy distributions for the positrons, electrons, and ions. By the end of the simulation ($3\,\mathrm{\mu s}$), the positrons have transferred significant energy into the electrons and ions.
@@
Figure 5: Mean kinetic energy of positrons (blue), electrons (red), and ions (green) as a function of time for the fiducial parameters. This simulation goes past the fiducial $3\,\mathrm{\mu s}$ up to $7.5\,\mathrm{\mu s}$ to show the increasing ion kinetic energies.
@@ 3.2 Suite of Simulations
To probe the range of non-linear effects in the simulations due to varying high-energy lepton fractions (from radioactive decay), we conduct a suite of simulations with different high-energy positron fractions (high-energy electrons are discussed for kilonovae in Section 4.5.1). This allows us to infer the magnetic field saturation levels corresponding to each positron fraction. For simplicity, the previously defined positron fraction $f_{p}$ can be generalized to the total decay lepton fraction.
@@ 4 Implications for Radioactive Transients
Here we consider how these plasma instabilities affect the late-time evolution of radioactively powered transients. First we outline the important decay chains and relevant derivations in Section 4.1. Thermonuclear/Type Ia SNe are considered in Section 4.2, Type II SNe in Section 4.3 and Stripped-Envelope supernovae (SE SNe) in Section 4.4, and we finish with kilonovae in Section 4.5.
@@ 4.1.1 Radioactive Decay in Transients
The luminosity of most SNe Ia and kilonovae (KNe) past $\sim 1$ day primarily results from the radioactive decay of isotopes produced during their explosion (e.g., Arnett, 1982; Li & Paczyński, 1998). In SNe, most of the synthesized mass comprises nuclear species with mass numbers ranging from $A\approx 12-70$, which are generally stable or proton-rich nuclides (e.g., Arnett, 1996; Seitenzahl et al., 2013). For these nuclei, fission and alpha decay are not significant. Radioactive decay in SNe progresses along an isobar towards more neutron-rich nuclides through two primary mechanisms: electron capture and positron emission. In electron capture, an inner atomic electron is captured by a nuclear proton, resulting in the emission of an electron neutrino. Positron emission, on the other hand, involves the decay of a nuclear proton into a neutron, accompanied by the emission of a positron and an electron neutrino. Both processes involve transitions to various nuclear levels in the daughter nucleus, with probabilities determined by statistical branching ratios (e.g., Arnett, 1996; Seitenzahl et al., 2013).
@@ 4.2.1 Overview and Physical Properties
SNe Ia are thermonuclear explosions of carbon/oxygen (C/O) white dwarfs (WDs; e.g., Hoyle & Fowler, 1960). Observational studies have measured the 56Ni mass ($M_{\mathrm{Ni}}$) in SNe Ia to range between $0.2$ and $1.6\,M_{\odot}$, with a median around $0.6\,M_{\odot}$ (e.g., Iwamoto et al., 1999; Seitenzahl et al., 2009; Scalzo et al., 2014; Childress et al., 2015; Scalzo et al., 2019) and the total ejecta mass ($M_{\mathrm{ej}}$) is between $0.8$ and $2.2\,M_{\odot}$, with a median around $1.4\,M_{\odot}$ (e.g., Scalzo et al., 2014, 2019).
@@
The isotopic composition of the ejecta is often stratified, with 56Ni predominantly concentrated near the center, while lighter elements are distributed toward the outer layers (e.g., Sasdelli et al., 2014; Ashall et al., 2014; Mazzali et al., 2015; Ashall et al., 2016). However, some models suggest that macroscopic mixing can result in a more extensive distribution of 56Ni throughout the ejecta (e.g., Seitenzahl et al., 2013; Piro & Morozova, 2016). For our analysis, we consider two limiting scenarios: one where 56Ni is centrally concentrated, with lighter elements confined to the outer layers, and another where 56Ni is homogeneously mixed throughout the ejecta. The first scenario aligns with expectations for normal SNe Ia, while the second is relevant to the low-luminosity Iax subclass of SNe Ia, thought to originate from incomplete deflagrations (e.g., Jha, 2017; Camacho-Neves et al., 2023). Although future work is needed to self-consistently incorporate plasma streaming instabilities into realistic explosion models, these simplified scenarios provide robust boundary conditions for exploring the range of possible outcomes.
@@ 4.2.2 Magnetic Fields
Figure 8a illustrates this evolution for SNe Ia, where the dark cyan band represents the range of possible magnetic field strengths for our adopted range of parameters (i.e., $M_{\mathrm{ej}}$, $M_{\mathrm{Ni}}$, $v_{\mathrm{ej}}$, $R_{0}$). The dashed line shows the magnetic field evolution for a set of typical SNe Ia parameters for the case of centrally concentrated 56Ni, and the dash-dotted line shows the evolution for 56Ni mixed within the ejecta. Both evolve similarly, except, centrally concentrated models have higher ion densities and thus a slightly higher magnetic field.
@@
With this magnetic field evolution, prompt synchrotron emission from the high-energy positrons is expected to peak in the $100\,\mathrm{MHz}$ to $1\,\mathrm{GHz}$ range, $\sim 20$ days after the explosion. However, for radio observatories such as LOFAR, SKAO and future radio facilities to be able to detect radio signatures from high-energy positrons, SNe Ia must be (1) Galactic or local-group SNe, and (2) in extremely clean environments with no circumstellar material. More details on synchrotron emission from SNe Ia positrons and its detectability are given in Appendix B.
@@ 4.3.1 Overview and Physical Properties
SNe II are the most common subclass of core-collapse supernovae that originate from the explosive death of massive stars ($M_{\mathrm{ZAMS}}>8\,\mathrm{M_{\odot}}$, e.g., Hoyle & Fowler, 1960; Woosley et al., 2002; Smartt, 2009). They are typically categorized by the presence of hydrogen in their spectra, distinguishing them from hydrogen-deficient SE SNe and thermonuclear SNe Ia. The light curves of SNe II are notably diverse, but often exhibit a plateau phase (Type IIP) that persists for several weeks, during which the luminosity remains nearly constant (e.g., Popov, 1993). This plateau results from the recombination of ionized hydrogen in the extended outer layers of the ejecta, effectively creating a diffusion barrier that slows the escape of photons. The plateau phase is followed by a decline as the supernova transitions to the radioactive decay phase, primarily powered by the decay of 56Ni to 56Co and subsequently to stable 56Fe (e.g., Arnett, 1982; Woosley & Weaver, 1986).
@@
Key observational properties, such as the nickel mass ($M_{\mathrm{Ni}}$) and ejecta mass ($M_{\mathrm{ej}}$), are typically inferred from the light curves of SNe II. The nickel mass can vary widely, with reported values ranging from $10^{-3}$ to $0.3\,\mathrm{M_{\odot}}$, and a median value of approximately $0.03\,\mathrm{M_{\odot}}$ (Hamuy, 2003; Müller et al., 2017; Anderson, 2019). The ejecta mass is typically $10\,\mathrm{M_{\odot}}$ with a range of $8$ to $30\,\mathrm{M_{\odot}}$ (e.g., Martinez et al., 2022). Ejecta velocities ($v_{\mathrm{ej}}$) usually determined from spectral features like the Feii $\lambda 5169$ absorption line span a range from $1000$ to $9000\,\mathrm{km\,s^{-1}}$, with a median velocity of $4000\,\mathrm{km\,s^{-1}}$ (Hamuy, 2003; Gutiérrez et al., 2017). Velocities of outer layers of ejecta are much faster, estimated with H and He lines at a median of $7000\,\mathrm{km\,s^{-1}}$, with a range of $2000$ to $14{,}000\,\mathrm{km\,s^{-1}}$ (Gutiérrez et al., 2017).
@@ 4.3.2 Magnetic Fields
The magnetic field generated by plasma instabilities in the ejecta begins with a large spread before $\sim 1$ day due to the wide range of possible explosion parameters. While the ranges of plasma-generated and progenitor magnetic fields overlap initially, the plasma-generated field surpasses the progenitor surface field within $\sim 1$ day in our typical cases (dashed and dash-dotted lines in Figure 8b). This rapid increase in difference aligns with the timescale on which ion densities decrease enough to reduce positron-electron annihilation, thereby allowing the instability to sustain itself and amplify the magnetic field. As the ejecta continue to expand, the ion density decreases, and consequently, the instantaneous saturated magnetic field weakens. This decrease follows a power-law decline, as indicated by the slope of approximately $-1.5$ in the logarithmic plot in Figure 8b. The plasma-generated magnetic field remains dominant over the interstellar medium (ISM) magnetic field until roughly $10^{5}$ days post-explosion, after which the ISM field strength becomes comparable.

### v2

@@ 1 Introduction
SNe Ia, in particular, are ideal environments for understanding lepton propagation in radioactively-powered ejecta. Despite the observational necessity, the mechanism behind the continued positron confinement in SNe Ia is unclear. Previous studies have suggested that extremely strong pre-existing magnetic fields ($>10^{6}\,\mathrm{G}$) can explain local positron trapping at late times ($>300$ days; Milne et al., 1999; Penney & Hoeflich, 2014; Hristov et al., 2021), but implies initial white dwarf (WD) surface fields of $>10^{9}\,\mathrm{G}$, surpassing all observed field strengths ($10^{3}-10^{9}\,\mathrm{G}$; Schmidt et al., 2003; Külebi et al., 2009; Ferrario et al., 2015). Another option is the delayed deposition of positron energy, which is predicted to flatten the light curve (Kushnir & Waxman, 2020), but has yet to be observationally confirmed.
@@
Plasma streaming instabilities offer a promising solution for positron confinement without the need for strong pre-existing magnetic fields. Plasma streaming instabilities draw energy from the bulk motion of one species and channel it into amplifying electromagnetic perturbations over time, disrupting the plasma’s equilibrium and leading to strong and often turbulent fields. These instabilities play a crucial role in the dynamics of high-energy astrophysical phenomena, but are often overlooked in transient simulations. They are instrumental in processes ranging from electron scattering in the solar wind (Roberg-Clark et al., 2018; Cattell et al., 2021) and intracluster medium of galaxy clusters (Roberg-Clark et al., 2018) to forward shock of $\gamma$-ray bursts (Sironi & Spitkovsky, 2011; Golant et al., 2025). Among the most pertinent plasma streaming instabilities arising from a net electron or positron current are filamentation instabilities (e.g., Weibel, 1959; Bell, 1978, 2004; Bret, 2009; Reville & Bell, 2012; Caprioli & Spitkovsky, 2013; Gupta et al., 2021). Such instabilities can grow exponentially from the density perturbations in the anisotropic flow of particles even without an initial magnetic field and saturate when the electron or positron gyro-radius (Larmor radius) becomes comparable to their inertial length. They can efficiently trap energetic particles by generating local magnetic fields that slow down or halt particle diffusion (Riquelme & Spitkovsky, 2011; Schroer et al., 2021; Gupta et al., 2021).
@@
In this work, we study plasma streaming instabilities in radioactive transients using simulations, taking SNe Ia as fiducial examples. Using fully kinetic particle-in-cell (PIC) simulations, we demonstrate how these instabilities arise and evolve in the presence of high-energy leptons, leading to self-generated magnetic fields that confine these particles even in the absence of an initial magnetic field. In Section 2, we establish the theoretical groundwork of plasma instabilities generated by high-energy leptons. In Section 3, we describe our PIC simulations and the results. In Section 4, we apply these findings to radioactive transients such as SNe Ia, CC SNe, and KNe. Finally, in Section 5, we summarize our conclusions.
@@ 3 Particle-in-Cell Simulations
(Spitkovsky, 2005; Hakobyan et al., 2023), a publicly available, fully parallelized PIC code specifically designed for astrophysical plasma simulations. The PIC method solves the coupled Vlasov-Maxwell system of equations for a collisionless plasma. Maxwell’s equations
@@ 3.1.2 Simulation Results
From this simple simulation with an initial magnetic field of zero, the system generates a saturated magnetic field in microseconds perpendicular perpendicular to the bulk flow. Since positrons are constrained to move along magnetic field lines, the perpendicular fields suggest that the positrons will remain fixed at the same mass coordinates in a radially symmetric explosion. To move outward, they must diffuse, indicating that they would deposit their energy locally instead.
@@
To examine how the particles are affected by the growing magnetic field, we show the initial and final distributions of particles in the 2D velocity space of $v_{x}$ and $v_{y}$, in Figure 4a and b. Due to the symmetry of the problem, $v_{y}$ and $v_{z}$ have the same statistical properties. The contours show the distributions of positrons, electrons, and ions. The inner levels of the contour represent a higher density of particles in velocity-space.
@@
Figure 4: (a) The initial and (b) final ($t=540\,\mathrm{\omega_{pe}^{-1}}$) velocity distributions for the positrons, electrons, and ions for the fiducial SN Ia simulation. The velocity of $v_{x,y}=0$ is marked by the dashed gray lines, indicating the point at which the particles would be at rest. (c) The initial and (d) final ($t=540\,\mathrm{\omega_{pe}^{-1}}$) kinetic energy distributions for the positrons, electrons, and ions. By the end of the simulation ($3\,\mathrm{\mu s}$), the positrons have transferred significant energy into the electrons and ions.
@@
Figure 5: Mean kinetic energy of positrons, electrons, and ions as a function of time for the fiducial parameters. This simulation goes past the fiducial $3\,\mathrm{\mu s}$ up to $7.5\,\mathrm{\mu s}$ to show the increasing ion kinetic energies.
@@ 3.2 Suite of Simulations
To probe the range of non-linear effects in the simulations due to varying high-energy lepton fractions (from radioactive decay), we conduct a suite of simulations with different high-energy positron fractions (high-energy electrons are discussed for KNe in Section 4.5.1). This allows us to infer the magnetic field saturation levels corresponding to each positron fraction. For simplicity, the previously defined positron fraction $f_{p}$ can be generalized to the total decay lepton fraction.
@@ 4 Implications for Radioactive Transients
Here we consider how these plasma instabilities affect the late-time evolution of radioactively powered transients. First we outline the important decay chains and relevant derivations in Section 4.1. Thermonuclear/SNe Ia are considered in Section 4.2, Type II SNe in Section 4.3 and Stripped-Envelope supernovae (SE SNe) in Section 4.4, and we finish with KNe in Section 4.5.
@@ 4.1.1 Radioactive Decay in Transients
The luminosity of most SNe Ia and KNe past $\sim 1$ day primarily results from the radioactive decay of isotopes produced during their explosion (e.g., Arnett, 1982; Li & Paczyński, 1998). In SNe, most of the synthesized mass comprises nuclear species with mass numbers ranging from $A\approx 12-70$, which are generally stable or proton-rich nuclides (e.g., Arnett, 1996; Seitenzahl et al., 2013). For these nuclei, fission and alpha decay are not significant. Radioactive decay in SNe progresses along an isobar towards more neutron-rich nuclides through two primary mechanisms: electron capture and positron emission. In electron capture, an inner atomic electron is captured by a nuclear proton, resulting in the emission of an electron neutrino. Positron emission, on the other hand, involves the decay of a nuclear proton into a neutron, accompanied by the emission of a positron and an electron neutrino. Both processes involve transitions to various nuclear levels in the daughter nucleus, with probabilities determined by statistical branching ratios (e.g., Arnett, 1996; Seitenzahl et al., 2013).
@@ 4.2.1 Overview and Physical Properties
SNe Ia are thermonuclear explosions of carbon/oxygen (C/O) WDs (e.g., Hoyle & Fowler, 1960). Observational studies have measured the 56Ni mass ($M_{\mathrm{Ni}}$) in SNe Ia to range between $0.2$ and $1.6\,M_{\odot}$, with a median around $0.6\,M_{\odot}$ (e.g., Iwamoto et al., 1999; Seitenzahl et al., 2009; Scalzo et al., 2014; Childress et al., 2015; Scalzo et al., 2019) and the total ejecta mass ($M_{\mathrm{ej}}$) is between $0.8$ and $2.2\,M_{\odot}$, with a median around $1.4\,M_{\odot}$ (e.g., Scalzo et al., 2014, 2019).
@@
The isotopic composition of the ejecta is often stratified, with 56Ni predominantly concentrated near the center, while lighter elements are distributed toward the outer layers (e.g., Ashall et al., 2014, 2016; Sasdelli et al., 2014; Mazzali et al., 2015). However, some models suggest that macroscopic mixing can result in a more extensive distribution of 56Ni throughout the ejecta (e.g., Seitenzahl et al., 2013; Piro & Morozova, 2016). For our analysis, we consider two limiting scenarios: one where 56Ni is centrally concentrated, with lighter elements confined to the outer layers, and another where 56Ni is homogeneously mixed throughout the ejecta. The first scenario aligns with expectations for normal SNe Ia, which are the most common (e.g., Desai et al., 2024), while the second is relevant to the less-common, low-luminosity Iax subclass of SNe Ia, thought to originate from incomplete deflagrations (e.g., Jha, 2017; Camacho-Neves et al., 2023). Although future work is needed to self-consistently incorporate plasma streaming instabilities into realistic explosion models, these simplified scenarios provide robust boundary conditions for exploring the range of possible outcomes.
@@ 4.2.2 Magnetic Fields
Figure 8a illustrates this evolution for SNe Ia, where the dark cyan band represents the range of possible magnetic field strengths for our adopted range of parameters (i.e., $M_{\mathrm{ej}}$, $M_{\mathrm{Ni}}$, $v_{\mathrm{ej}}$, and $R_{0}$). The dashed line shows the magnetic field evolution for a set of typical SNe Ia parameters for the case of centrally concentrated 56Ni, and the dash-dotted line shows the evolution for 56Ni mixed within the ejecta. Both evolve similarly, except, centrally concentrated models have higher ion densities and thus a slightly higher magnetic field.
@@
With this magnetic field evolution, prompt synchrotron emission from the high-energy positrons is expected to peak in the $100\,\mathrm{MHz}$ to $1\,\mathrm{GHz}$ range, $\sim 20$ days after the explosion. However, for radio observatories such as LOw Frequency ARray (LOFAR), Square Kilometre Array Observatory (SKAO) and future radio facilities to be able to detect radio signatures from high-energy positrons, SNe Ia must be (1) Galactic or local-group SNe, and (2) in extremely clean environments with no CSM. More details on synchrotron emission from SNe Ia positrons and its detectability are given in Appendix B.
@@ 4.3.1 Overview and Physical Properties
SNe II are the most common subclass of CC SNe that originate from the explosive death of massive stars ($M_{\mathrm{ZAMS}}>8\,\mathrm{M_{\odot}}$, e.g., Hoyle & Fowler, 1960; Woosley et al., 2002; Smartt, 2009). They are typically categorized by the presence of hydrogen in their spectra, distinguishing them from hydrogen-deficient SE SNe and thermonuclear SNe Ia. The light curves of SNe II are notably diverse, but often exhibit a plateau phase (Type IIP) that persists for several weeks, during which the luminosity remains nearly constant (e.g., Popov, 1993). This plateau results from the recombination of ionized hydrogen in the extended outer layers of the ejecta, effectively creating a diffusion barrier that slows the escape of photons. The plateau phase is followed by a decline as the supernova transitions to the radioactive decay phase, primarily powered by the decay of 56Ni to 56Co and subsequently to stable 56Fe (e.g., Arnett, 1982; Woosley & Weaver, 1986).
@@
Key observational properties, such as $M_{\mathrm{Ni}}$ and $M_{\mathrm{ej}}$, are typically inferred from the light curves of SNe II. The nickel mass can vary widely, with reported values ranging from $10^{-3}$ to $0.3\,\mathrm{M_{\odot}}$, and a median value of approximately $0.03\,\mathrm{M_{\odot}}$ (Hamuy, 2003; Müller et al., 2017; Anderson, 2019; Rodríguez et al., 2021). The ejecta mass is typically $10\,\mathrm{M_{\odot}}$ with a range of $8$ to $30\,\mathrm{M_{\odot}}$ (e.g., Martinez et al., 2022). Ejecta velocities ($v_{\mathrm{ej}}$) usually determined from spectral features like the Feii $\lambda 5169$ absorption line span a range from $1000$ to $9000\,\mathrm{km\,s^{-1}}$, with a median velocity of $4000\,\mathrm{km\,s^{-1}}$ (Hamuy, 2003; Gutiérrez et al., 2017). Velocities of outer layers of ejecta are much faster, estimated with H and He lines at a median of $7000\,\mathrm{km\,s^{-1}}$, with a range of $2000$ to $14{,}000\,\mathrm{km\,s^{-1}}$ (Gutiérrez et al., 2017).
@@ 4.3.2 Magnetic Fields
The magnetic field generated by plasma instabilities in the ejecta begins with a large spread before $\sim 1$ day due to the wide range of possible explosion parameters. While the ranges of plasma-generated and progenitor magnetic fields overlap initially, the plasma-generated field surpasses the progenitor surface field within $\sim 1$ day in our typical cases (dashed and dash-dotted lines in Figure 8b). This rapid increase in difference aligns with the timescale on which ion densities decrease enough to reduce positron-electron annihilation, thereby allowing the instability to sustain itself and amplify the magnetic field. As the ejecta continue to expand, the ion density decreases, and consequently, the instantaneous saturated magnetic field weakens. This decrease follows a power-law decline, as indicated by the slope of approximately $-1.5$ in the logarithmic plot in Figure 8b. The plasma-generated magnetic field remains dominant over the ISM magnetic field until roughly $10^{5}$ days post-explosion, after which the ISM field strength becomes comparable.